OpenStudy (anonymous):

State the equation of the axis of symmetry for -(x+2)(x-2)

OpenStudy (anonymous):

-(x+2)(X-2)=-(x^2-4)=-x^2+4 So the equation of the axis of symmetry for the equation you asked is X=0.

OpenStudy (anonymous):

You can remember that it is a good way to make any equation like you asked into the form of Ax^2+Bx+C. So the equation of the axis is -B/2A

OpenStudy (anonymous):

$$y=-x^2+4 \Rightarrow (y-4)=-(x-0)^2 $$ the axis of symmetry is the y axis which equation is x=0